OpenStudy (anonymous):

Solve for b. d = 3a + 3b

OpenStudy (anonymous):

d = 3a + 3b --- subtract 3a from both sides d - 3a = 3b -- divide both sides by 3 (d - 3a) / 3 = b or d/3 - a = b

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

anytime :)